A student has to select 2 books from a collection 25 biology, 21 mathematics,
and 14 chemistry books. How many choices are possible if: both books are to
be on the same subject?

Answer by Edwin McCravy(20054)

25 biology books Choose 2        <-- 25C2 ways
      or                             plus
21 mathematics books Choose 2    <-- 21C2 ways
      or                             plus 
14 chemistry books Choose 2      <-- 14C2 ways
      
25C2 + 21C2 + 14C2 = 300 + 210 + 91 = 601 ways.

Edwin